OpenStudy (anonymous):

√5x-1 = √4x+9 *In both problems all of the numbers and letters are in the radical

OpenStudy (anonymous):

Square both sides.$$\sqrt{5x-1}=\sqrt{4x+9}\\5x-1=4x+9\\5x-4x=9+1\\x=10$$

OpenStudy (anonymous):

@oldrin.bataku thank you